# The Effect of Sampling Temperature on Problem Solving in Large Language Models

### v1

@@ 4.1 Limitations
There were several limitations to our research study:
First, our study was limited to a small subset of problems, problem domains, and problem-solving tasks. As a result, our findings may not hold for larger data sets, different problem domains, or other types of problem-solving tasks.
Second, we only explored a small subset of prompt-engineering techniques using a single prompt-and-response cycle with one-shot in-context learning. As a result, there may be more complex prompts or agent architectures that better leverage sampling temperature in their problem-solving capabilities.
Third, due to cost considerations, we could only fully explore the landscape of system prompts and problem domains with GPT-3.5. As a result, there may exist other combinations of system prompts and problem domains where other LLMs may better utilize sampling temperature for problem-solving.
@@ 4.2 Implications
This research study provides empirical evidence that changes to sampling temperature in the range of 0.0 to 1.0 do not affect the problem-solving capabilities of LLMs on MCQA problems.
Answering this question could save AI engineers significant time and resources evaluating LLM solutions with various sampling temperatures. In addition, it may reduce unproductive debates in the prompt-engineering community about optimal sampling temperatures for problem-solving tasks.
This research also provides more general insight for AI researchers studying model hallucination and state-space search with LLMs. Our results show that increasing LLM temperature up to 1.0 does not create hallucinations that lead to incorrect MCQA solutions. In addition, higher temperatures do not appear to improve MCQA solution-space search in ways that lead to correct solutions more often.
@@ 4.3 Future Research
First, we should perform additional experiments with more MCQA problems and problem domains. We recommend specifically targeting tasks and problem domains that require more creative solutions or lateral “out-of-the-box” thinking.
@@
Third, we recommend performing these experiments with additional LLMs. Other proprietary and open-source LLMs may utilize temperature in ways that benefit their specific models but did not benefit GPT-3.5 and GPT-4.
@@ 5 Conclusion
This research study provided an empirical investigation of the effect of sampling temperature on the problem-solving performance of LLMs across multiple problem domains.
We demonstrated that changes in sampling temperature within the range of 0.0 to 1.0 do not produce statistically significant differences in problem-solving performance on MCQA problems across multiple LLMs, prompt-engineering techniques, and problem domains.
These results have practical implications for AI engineers using LLMs to create new AI systems. In addition, they have theoretical implications for AI researchers studying model hallucination and solution-space search with LLMs.
Impact Statement
This paper presents work whose goal is to advance the field of Artificial Intelligence and Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically addressed here.
Appendix A Prompts
@@
  B: Annapolis is a the capital of the State of Maryland.
@@
  You are correct that Johns Hopkins is located in the State of Baltimore.

### v2

@@ 4 Discussion
4.1 Interpretation
Based on these results, changes in temperature from 0.0 to 1.0 do not have a statistically significant effect on the problem-solving performance of LLMs. These results appear to generalize across LLMs, prompt-engineering techniques, and problem domains. However, there were a few potential exceptions to these general findings.
Therefore, we recommend setting an LLM’s sampling temperature to 0.0 for problem-solving tasks. This temperature maximizes reproducibility without compromising accuracy. In addition, it helps avoid the inevitable dropoff in performance that occurs beyond a temperature of 1.0. However, exceptions to this guidance should be taken into consideration.
@@ 4.2 Limitations
There were several limitations to our research study due to cost and runtime considerations:
First, our study was limited to a subset of popular LLMs. As a result, our findings may not hold for other LLMs that were excluded from our experiment.
Second, we only explored a subset of prompt-engineering techniques using a single prompt-and-response cycle with one-shot in-context learning. As a result, there may be more complex prompts or agent architectures that better leverage sampling temperature for creativity in their problem-solving capabilities.
Third, our study was limited to a subset of problems, problem domains, and problem-solving tasks. As a result, our findings may not hold for larger data sets, different problem domains, or other types of problem-solving tasks.
Fourth, we had to limit the sampling temperature we explored from 0.0 to 1.0 for all combinations of models, prompts, and exams, except for GPT-3.5 using CoT prompting on the 100-question exam. As a result, the temperature hyperparameter of other LLMs may operate differently at temperatures above 1.0.
Finally, we could only explore a subset of the various combinations of models, prompts, exams, and temperatures. As a result, other combinations of LLMs, prompt-engineering techniques, and problem domains may exist where temperature plays a more important role in problem-solving performance.
@@ 4.3 Implications
This research study provides empirical evidence that changes in sampling temperature in the range of 0.0 to 1.0 do not impact the problem-solving capabilities of LLMs on MCQA problems.
Answering this question may save AI engineers significant time and resources evaluating various sampling temperatures for their LLM agents and applications. In addition, it may reduce unproductive debates in the prompt-engineering community regarding the optimal sampling temperatures for various problem-solving tasks.
This research also provides broader insights for AI researchers studying model hallucination and problem-solution state-space search with LLMs. Our results show that increasing LLM temperature up to 1.0 does not cause the LLM to hallucinate in ways that lead to incorrect MCQA solutions. In addition, higher temperatures do not appear to improve MCQA solution-space search in ways that lead to correct solutions more often.
@@ 4.4 Future Research
First, we recommend conducting this experiment with additional LLMs. Other proprietary and open-source LLMs may utilize temperature in ways that benefit their specific models but did not benefit the LLMs we tested.
@@
Third, we recommend conducting additional experiments with more MCQA problems and problem domains. We recommend specifically targeting tasks and problem domains that require more creative solutions or lateral “out-of-the-box” thinking.
Fourth, we recommend extending the sampling temperature range until accuracy drops to zero for each LLM, prompt, and exam. However, it should be noted that as the generated text becomes more random, the number of tokens in each response increases significantly, leading to a considerable increase in runtime and cost to explore temperatures above 1.0.
@@ 5 Conclusion
This research study empirically investigated the effect of sampling temperature on the problem-solving performance of LLMs across multiple problem domains.
We demonstrated that changes in sampling temperature from 0.0 to 1.0 do not produce statistically significant differences in problem-solving performance on MCQA problems across multiple LLMs, prompt-engineering techniques, and problem domains.
These results have practical implications for AI engineers using LLMs to develop new AI systems. Additionally, they have theoretical implications for AI researchers studying model hallucination and solution-space search with LLMs.
Appendix A Appendix
@@
  B: Annapolis is the capital of the State of Maryland.
@@
  You are correct that Johns Hopkins is located in the State of Maryland.